OpenStudy (daenio):

There are 15 students in our class that like root beer, 23 students that like cola and 8 students that like both root beer and cola. If there are 50 students in our class, what is the probability of choosing a random student such that this student likes only one type of pop? Question 3 options: a) 92% b) 60% c) 76% d) 44%

OpenStudy (daenio):

@kropot72

OpenStudy (daenio):

@Mertsj @aroub @ZhangYan

OpenStudy (kropot72):

The number of students that like one type is 15 + 23 = 38 The required probability is\[\frac{38}{50}\times 100=\] Can you finish it?

OpenStudy (daenio):

76% and 60% are not the correct solutions.

OpenStudy (daenio):

@Zarkon @imranmeah91

OpenStudy (kaederfds):

Just like cola = 15 just like root beer = 7 22/50.100% = 44% ^^

OpenStudy (daenio):

Thanks!